1.3 The equation mv gl sin 0 +kl2 describes the energy of an object with mass m amd initial speed vo sliding a distance l up a frictionless ramp at inclination angle θ, all the while compressing a spring with stiffness k as it comes to rest. Solve the equation for I, the distance the object moves up the incline and compresses the spring. Please be as detailed as possible so I can follow each step to learn this. Thanks!
Answer #1

Solution :

We are given that

malsn........................(1)

Above equation represents the conservation of mechanical energy and we have to find "l" from this equation.

If we look carefully then we can see that this equation is a quadratic equation in "l".

First, let's rearrange this equation in standard form i.e. in the below form

ax^2+bx+c=0

rearranging equation (1),we have

kl^2+2mglsin heta-mv^2_o=0..................(2)

Now solution of a quadratic equation is given by

ас CL............(3)

comparing equation (1) and (2)

r = l, a = k, b = 2nngsind, c = -?mi

Now to find "l" put all values in equation (3)

l= rac{-2mgsin hetapm sqrt{4m^2g^2sin^2 heta+4kmv^2_o}}{2k}

So their are two possible values of "l" but one of them is negative so the only possible value of "l" is

l= rac{-2mgsin heta+ sqrt{4m^2g^2sin^2 heta+4kmv^2_o}}{2k}......................Answer.